# Latent Assistance Networks: Rediscovering Hyperbolic Tangents in RL

### v1

@@ C.1 Shallow and Deep Function Approximation
To further showcase the effect of activations on complex function approximation, we compare the single hidden layer Tanh (LaN) network from Fig. 2 with a deep ReLU and Tanh network containing three hidden layers each. The comparison with shallow networks can be found in Fig. 12(a) and a comparison with deep networks can be found in Fig. 12(b).
@@
(a) Comparison of shallow networks for a nonlinear regression task. The Tanh and ReLU networks have a single hidden layer of 200 neurons, while the Tanh (LaN) has a single hidden layer of 100 neurons but two preceding linear layers. The Tanh and ReLU networks have 601 parameters, while the Tanh (LaN) network has 501 parameters. As found by Gulcehre et al. (2022), a shallow network activated by ReLU has a lower effective rank and consequently reduced network expressivity as compared to a Tanh activated network. Using a LaN, we achieve better function approximation while using less parameters.
@@
(b) Comparison of two deep networks and one shallow network for the same nonlinear regression task. The Tanh and ReLU networks have 3 hidden layers of 200 neurons each, while the Tanh (LaN) network remains shallow. In line with common observations in deep learning, the ReLU activation thrives in deeper networks, in contrast to the Tanh activation. Interestingly, the shallow Tanh (LaN) network still achieves better function approximation with only 0.6% of the deeper networks’ parameters (81001 vs 501). No hyperparameter tuning or architecture search has been applied. Additional tests using deep Tanh (LaN) networks gave similar function approximation as compared to the shallow Tanh (LaN) network.
C.2 Validation of dying neuron probability derivations
@@
As discussed in Section 3, the effect of adding a Latent assistance Network (LaN) to the original representation strongly depends on the activation function. These derivations are empirically validated by the results in Fig. 13. In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1. This could explain the very slight increase in dead neurons when using a LaN with a sigmoid.
(a) Sigmoid with and without LaN
(b) Tanh with and without LaN
(c) ReLU with and without LaN
@@
Figure 13: By evaluating the effect of a LaN on dying neurons through the lens of probability theory, we predicted that only the hyperbolic tangent benefits in this metric. Specifically, only a hyperbolic tangent was speculated to have a decrease in dying neurons. Using a LaN on a sigmoid activation would have no notable difference, and for the ReLU activation we expected an increase in dead neurons. This empirically validates our theory in Section 3.
C.3 Combining ReLU with a LaN
@@
Additional Atari experiments are provided comparing a ReLU activation with and without a LaN. The results can be seen in fig. 14.
@@
Figure 14: As a Rectified Linear Unit creates sparse representations, it does not benefit from using a LaN, since the final representation will consist of the element-wise product between two sparse representations. Therefore, a decrease in both performance and effective rank and an increase in dead neurons can be expected.
@@ D.1 Evaluation Details
Figure 15: Human-Normalized performance (in multiples) with the standard deviation over the means in the Atari domain for 10M iterations (40M Frames)
D.2 Individual Environments
Figure 16: DQN Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. To avoid cherry-picking, no other Atari games were tested.
Figure 17: PPO Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. To avoid cherry-picking, no other Atari games were tested.

### v2

@@ C.1 Shallow and Deep Function Approximation
To further showcase the effect of activations on complex function approximation, we compare the single hidden layer Tanh (HR) network from Fig. 2 with a deep ReLU and Tanh network containing three hidden layers each. The comparison with shallow networks can be found in Fig. 13(a) and a comparison with deep networks can be found in Fig. 13(b).
@@
(a) Comparison of shallow networks for a nonlinear regression task. The Tanh and ReLU networks have a single hidden layer of 200 neurons, while the Tanh (HR) has a single hidden layer of 100 neurons but two preceding linear layers. The Tanh and ReLU networks have 601 parameters, while the Tanh (HR) network has 501 parameters. As found by Gulcehre et al. (2022), a shallow network activated by ReLU has a lower effective rank and consequently reduced network expressivity as compared to a Tanh activated network. Using a Hadamard representation, we achieve better function approximation while using less parameters.
@@
(b) Comparison of two deep networks and one shallow network for the same nonlinear regression task. The Tanh and ReLU networks have 3 hidden layers of 200 neurons each, while the Tanh (HR) network remains shallow. In line with common observations in deep learning, the ReLU activation thrives in deeper networks, in contrast to the Tanh activation. Interestingly, the shallow Tanh (HR) network still achieves better function approximation with only 0.6% of the deeper networks’ parameters (81001 vs 501). No hyperparameter tuning or architecture search has been applied. Additional tests using deep Tanh (HR) networks gave similar function approximation as compared to the shallow Tanh (HR) network.
C.2 Increasing Representation Parameters
(a) Performance
(b) Dead Neurons
(c) Effective Rank
Figure 14: Comparison of a normal hyperbolic tangent (Tanh), a hyperbolic tangent with a higher representation dimension $z_{t}\in\mathbb{R}^{512\rightarrow 1024}$ and a Hadamard representation using hyperbolic tangents. Comparisons are done on Performance (a), the fraction of dead neurons (b) and the effective rank of the representation $z_{t}$ (c). Naturally, increasing the representation dimension $z_{t}$ increases the effective rank of the representation, but using a larger representation dimension is not always preferable as it can negatively impact actual convergence.
C.3 Validating dying neuron probability derivations
@@
As discussed in Section 3, the effect of using a Hadamard representation strongly depends on the activation function. These derivations are empirically validated by the results in Fig. 15. In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1. This could explain the very slight increase in dead neurons when using an HR with activations.
(a) Sigmoid with and without HR
(b) Tanh with and without HR
(c) ReLU with and without HR
@@
Figure 15: By evaluating the effect of an HR on dying neurons through the lens of probability theory, we predicted that only the hyperbolic tangent benefits in this metric. Specifically, only a hyperbolic tangent was speculated to have a decrease in dying neurons. Using an HR with sigmoid activations would have no notable difference, and for an HR with ReLU activations an increase in dead neurons was expected. This empirically validates our hypotheses in Section 3.
C.4 ReLU activated Hadamard representation
@@
Additional Atari experiments are provided comparing a ReLU activation with and without an HR. The normalized scores, dying neurons and the effective rank during training can be seen in fig. 16.
@@
Figure 16: As a Rectified Linear Unit creates sparse representations, it does not benefit from using an HR, since the final representation will consist of the Hadamard product between two sparse representations. Therefore, a decrease in both performance and effective rank and an increase in dead neurons can be expected.
@@ D.1 Evaluation Details
Figure 17: Human-Normalized performance (in multiples) with the standard deviation over the means in the Atari domain for 10M iterations (40M Frames).
D.2 Individual Environment Scores
Figure 18: DQN Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. No other games were tested.
Figure 19: PPO Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. No other games were tested.